# Linear Preference Optimization: Decoupled Gradient Control via Absolute Regularization

### v1

@@ 3.3 Preference Pairs Construction
In the SPIN Chen et al. (2024), it is indicated that following Supervised Fine Tuning (SFT), a general model’s output still exhibits certain discrepancies when contrasted with the Ground Truth. Iterative DPO Pang et al. (2024) mentions that when constructing preference pairs, a Reward Model can be employed to select samples with the highest and lowest scores to form these pairs. Consequently, we propose a novel method for constructing preference pairs without the involvement of a reward model: the chosen sample is deemed a sufficiently good answer, whereas the reject sample is generated by the SFT model inference hyperparameters where both top-p and temperature are set to 1.0. The specific algorithm pseudocode is shown in Algorithm 1.
@@ 4 Experiments
In this section, we first introduce the experimental setup and then present comparative experimental results of the models. To validate the effectiveness of the algorithm, comprehensive experiments were conducted in three domains: general text tasks, domain-specific mathematical text tasks, and TTS speech generation tasks.
@@ 4.1 Results on General Tasks
We selected Qwen2.5Team (2024) as the base model. Infinity-InstructLi et al. (2025) is an open-source, high-quality dataset. For supervised fine-tuning (SFT), we randomly selected 290k examples from the Infinity-Instruct dataset. The resulting model after SFT is denoted as qwen2.5-SFT. Detailed SFT experimental settings and the training procedure are provided in the Appendix B.
@@
1) Infinity-Preference: An open-source, high-quality preference dataset characterized by subtle distinctions between chosen and rejected responses, minimal noise, and greater learning difficulty.
2) Infinity-instruct-1w: We randomly selected 10,000 samples from the remaining Infinity-instruct data. The responses of the original Infinity-instruct dataset serve as the chosen set, while data generated by Qwen2.5-SFT (under temperature=1.0, top_p=1.0) serves as the rejected set. This constructed training dataset is of lower quality than Infinity-Preference.
Previous research has demonstrated that GPT-4 evaluations closely align with human assessments and are significantly more cost-effectiveZheng et al. (2023). Thus, we select two GPT-4-based evaluation benchmarks: MT-benchBai et al. (2024) and AlignbenchLiu et al. (2023). MT-bench covers diverse categories including writing, STEM, role-play, reasoning, mathematics, humanities, information extraction, and coding, while also evaluating multi-turn dialogue performance. We employ GPT-4 as the judge model within this benchmark. Similarly, Alignbench includes evaluation categories such as mathematical computation, role-playing, logical reasoning, and text writing.
We employ vanilla DPORafailov et al. (2023) as our baseline comparison algorithm. For DPO experiments, we follow the official setup, setting the $\beta$ hyperparameter to 0.1. The experimental setup for LPO are presented in the Appendix Table 9.
For comprehensive tasks, we conducted detailed experiments on both MT-bench and Alignbench. MT-bench encompasses diverse categories including writing, STEM, role-playing, reasoning, mathematics, humanities, information extraction, and coding. Similarly, Alignbench covers various dimensions such as mathematical computation, role-playing, logical reasoning, and text composition. In both benchmarks, we employ GPT-4 as the evaluation judge model.
@@
Based on Table 1 and Table 2 and Table 3 and Table 4, it is observed that the LPO algorithm achieves significant improvements over the SFT model: a 6.37% gain on MT-bench and a 2.24% improvement on Alignbench when trained on the Infinity-Preference dataset, along with a 4.81% enhancement on MT-bench when trained on the Infinity-Instruct-1w dataset. While the model trained with DPO on the Infinity-Preference dataset showed notable gains, its performance on the Infinity-Instruct-1w dataset exhibited a slight decline on MT-bench and a more pronounced drop on Alignbench. Further analysis reveals that paired data constructed from Infinity-Preference exhibits finer distinctions and presents greater learning challenges, whereas pairs derived from Infinity-Instruct-1w display more pronounced differentiability and lower learning difficulty. These findings indicate that the LPO algorithm demonstrates superior robustness, delivering consistent gains across diverse datasets, while DPO exhibits higher sensitivity to training data and appears prone to overfitting on simpler datasets. LPO demonstrates substantial gains in logical reasoning, while DPO shows more significant improvements in scenarios like Q&A. We have further assessed the model’s effectiveness specifically within the mathematics vertical.
@@ 4.2 Results on Math Tasks
We initialize our model using the SFT model pre-trained on general tasks. During the alignment phase, we train the model using the dataset constructed via step-dpoLai et al. (2024). Detailed experimental configurations are provided in the Appendix C.
We conducted extensive tests on the GSM8KCobbe et al. (2021) benchmark. To better reflect real-world usage scenarios, we employed a zero-shot approach during inference. Here, ”Qwen2.5-Instruct” denotes the officially released model from the Qwen team. The experimental results are presented in Table 5.
@@ 4.3 Results on Text-to-Speech Tasks
The LPO training dataset is constructed as follows: Using each sample’s prompt as input, we generate three candidate responses. These candidates are then paired with the sample’s reference answer to form three preference pairs, comprising the final training data. Corresponding training parameters are detailed in Appendix C.
@@ 4.4 Results on Automatic Speech Recognition Task
Automatic Speech Recognition (ASR) refers to the process of transforming spoken words into textual form, a task that involves detecting words within audio inputs and converting them into written language. The objective is to achieve a precise conversion of speech into text. To assess the LPO’s performance on ASR, we conducted sequential SFT and LPO training on the expanded Qwen-2.5-7B model, leveraging both AISHELL-1 (Chinese) and LibriSpeech (English) corpora.
The LPO training data were prepared using two different methodologies: (i) Model-based: Inspired by TTS data synthesis methodologies, we leveraged the SFT model to generate LPO candidate samples and then paired with reference samples. (ii) Perturbation-based: Text perturbation strategies Li et al. (2022) were adopted to generate candidate variants via systematic corruption of reference sentences using insertion, deletion and repetition operators at controlled noise ratios ($\eta=0.1$). Candidate samples from these two construction methods exhibit divergent properties: the model-based method primarily generates homophonic heterographs (phonetically identical or similar but semantically distinct words), while the perturbation-based method produces samples with quantifiable stochasticity.
@@
AISHELL-1 Model-based 1.0 3.513 3.868
2.0 3.517
3.0 3.526
Perturbation-based 1.0 3.571
2.0 3.332
3.0 3.554
LibriSpeech-test-clean Model-based 1.0 7.117 7.222
2.0 7.065
3.0 6.919
Perturbation-based 1.0 6.992
2.0 7.182
3.0 10.904
@@ 4.5 Analysis of multi-epoch with different $r_{2}$
For the general tasks, following the experimental setup detailed in Section 4.1, $r_{2}$ coefficients were selected from the range [1, 1.5, 2, 3]. For vertically specific math tasks, using the configuration described in Section 4.2, $r_{2}$ coefficients were tested at values of [0.1, 0.2, 1, 2]. The corresponding experimental results are illustrated in Figure 3(a) and Figure 3(b).
@@
As shown in Figures 3(a) and 3(b), our experimental results demonstrate two key conclusions: 1) Model performance varies on the leaderboard with different r2 coefficients. Thus, adjusting the r2 coefficient is necessary to prevent a rapid decline in rejection rate that causes overfitting. 2) The difficulty of learning varies across tasks. The r2 coefficient should be adjusted based on the rate of loss decrease for different tasks.

### v2

@@ 3.3 Preference Pairs Construction
In the SPIN Chen et al. (2024), it is noted that after SFT training, a general model’s output still displays certain discrepancies when compared to the Ground Truth. Iterative DPO Pang et al. (2024) suggests that a reward model can be utilized to select samples with the highest and lowest scores for constructing preference pairs.
Consequently, we propose a novel method for constructing preference pairs without relying on a reward model. In this approach, the chosen sample is considered a sufficiently good answer, while the rejected sample is generated using the SFT model’s inference hyperparameters, with both top-p and temperature set to 1.0.
The specific algorithm for this preference pair construction is detailed in Algorithm 1.
This method allows for the effective generation of preference pairs while reducing dependency on additional models, enhancing efficiency in the optimization process.
@@ 4 Experiments
In this section, we first introduce the experimental setup and subsequently present the comparative experimental results of the models. To validate the effectiveness of the proposed algorithm, comprehensive experiments were conducted across three distinct domains: general text tasks, domain-specific mathematical text tasks, text-to-speech (TTS) speech generation tasks, and automatic speech recognition (ASR) tasks.
The experimental framework involved selecting a diverse set of benchmarks to assess the performance of the models effectively. Each domain was meticulously designed to evaluate specific capabilities:
  1. 1.
    General Text Tasks: These tasks focused on a variety of language understanding and generation challenges, including writing, summarization, and question answering.
  2. 2.
    Domain-Specific Mathematical Text Tasks: This segment aimed to test the models on mathematical reasoning and problem-solving capabilities, utilizing datasets specifically curated for this purpose.
  3. 3.
    TTS Speech Generation Tasks: In this area, models were evaluated on their ability to synthesize natural-sounding speech, focusing on fidelity, stability, and emotional expressiveness.
  4. 4.
    ASR Speech Recognition Tasks: In this domain, models were assessed on their proficiency in converting spoken language into accurate and robust textual transcriptions, with an emphasis on handling diverse accents, background noise, and varying speech rates.
The results will be presented in terms of performance metrics relevant to each task type, showcasing the improvements achieved through the application of the proposed method compared to baseline models.
@@ 4.1 Results on General Tasks
We selected Qwen2.5 Team (2024) as the base model for our experiments. Infinity-Instruct Li et al. (2025) serves as an open-source, high-quality dataset. For SFT phase, we randomly selected 290k examples from the Infinity-Instruct dataset. The model resulting from this SFT process is denoted as qwen2.5-SFT. Detailed configurations and the training procedure for the SFT are provided in the Appendix B.
@@
1) Infinity-Preference: This is an open-source, high-quality preference dataset characterized by subtle distinctions between chosen and rejected responses. It features minimal noise and presents greater learning difficulty, making it an ideal candidate for robust evaluation.
2) Infinity-instruct-1w: For this dataset, we randomly selected 10,000 samples from the remaining Infinity-Instruct data. The responses from the original Infinity-Instruct dataset are designated as the chosen set, while the rejected set comprises data generated by Qwen2.5-SFT, utilizing temperature and top-p parameters set to 1.0. It is important to note that this constructed training dataset is of lower quality compared to Infinity-Preference.
Previous research has shown that GPT-4 evaluations closely align with human assessments and are significantly more cost-effective Zheng et al. (2023). Consequently, we selected two GPT-4-based evaluation benchmarks: MT-bench Bai et al. (2024) and Alignbench Liu et al. (2023).
MT-bench encompasses a variety of categories, including writing, STEM, role-play, reasoning, mathematics, humanities, information extraction, and coding, while also assessing multi-turn dialogue performance. For this benchmark, we employ GPT-4 as the judge model. In a similar vein, Alignbench includes evaluation categories such as mathematical computation, role-playing, logical reasoning, and text writing.
For comprehensive tasks, we conducted detailed experiments on both MT-bench and Alignbench. In both benchmarks, we employ GPT-4 as the evaluation judge model.
We employ vanilla DPO Rafailov et al. (2023) as our baseline comparison algorithm. For the DPO experiments, we adhere to the official setup, configuring the $\beta$ hyperparameter to 0.1. The experimental setup for LPO is detailed in Appendix Table 9.
@@
Shown in Table 1, 2, 3 and 4, the LPO algorithm achieves significant improvements over the SFT model: a 6.37% gain on MT-bench and a 2.24% improvement on Alignbench when trained on the Infinity-Preference dataset. Additionally, there is a 4.81% enhancement on MT-bench when using the Infinity-Instruct-1w dataset.
While the model trained with DPO on the Infinity-Preference dataset showed notable gains, its performance on the Infinity-Instruct-1w dataset exhibited a slight decline on MT-bench and a more pronounced drop on Alignbench. Further analysis reveals that paired data constructed from Infinity-Preference exhibits finer distinctions and presents greater learning challenges. In contrast, pairs derived from Infinity-Instruct-1w display more pronounced differentiability and lower learning difficulty.
These findings indicate that the LPO algorithm demonstrates superior robustness, delivering consistent gains across diverse datasets. In contrast, DPO exhibits higher sensitivity to training data and appears prone to overfitting on simpler datasets. LPO demonstrates substantial gains in logical reasoning, while DPO shows more significant improvements in scenarios like Q&A. We have further assessed the model’s effectiveness specifically within the mathematics vertical.
@@ 4.2 Results on Math Tasks
We initialize our model using the SFT model that has been pre-trained on general tasks. During the alignment phase, we train the model using the dataset constructed via step-DPO Lai et al. (2024). Detailed experimental configurations are provided in the Appendix C.
We conducted extensive tests on the GSM8K Cobbe et al. (2021) benchmark. To better reflect real-world usage scenarios, we employed a zero-shot approach during inference. Here, “Qwen2.5-Instruct” denotes the officially released model from the Qwen team. The experimental results are presented in Table 5.
@@ 4.3 Results on Text-to-Speech Tasks
The LPO training dataset is constructed as follows: Using each sample’s prompt as input, we generate three candidate responses. These candidates are then paired with the sample’s reference answer to form three preference pairs, which comprise the final training data. The corresponding training parameters are detailed in Appendix C.
@@ 4.4 Results on Automatic Speech Recognition Task
Automatic Speech Recognition (ASR) refers to the process of transforming spoken words into textual form, a task that involves detecting words within audio inputs and converting them into written language. The objective is to achieve a precise conversion of speech into text.
To assess the LPO’s performance on ASR, we conducted sequential SFT and LPO training on the expanded Qwen-2.5-7B model, leveraging both AISHELL-1 (Chinese) and LibriSpeech (English) corpora.
The LPO training data were prepared using two different methodologies:
  • Model-based: Inspired by TTS data synthesis methodologies, we leveraged the SFT model to generate LPO reject samples, which were then paired with reference samples.
  • Perturbation-based: Text perturbation strategies Li et al. (2022) were adopted to create reject samples through systematic corruption of reference sentences, employing insertion, deletion, and repetition operators at controlled noise ratios ($\eta=0.1$).
Reject samples derived from these two construction methods exhibit divergent properties: the model-based method primarily generates homophonic heterographs (phonetically identical or similar but semantically distinct words), while the perturbation-based method produces samples with quantifiable stochasticity.
Note that for both model-based and perturbation-based methods, it is necessary to exclude data pairs where the chosen and reject samples are the same. Furthermore, data quality has a far greater impact on model performance than data quantity. In our experiments, we found that 1k high-quality samples contributed more significantly to the model’s effectiveness and stability than 200k ordinary samples.
@@
AISHELL-1 Model-based 1.0 3.583 3.868
2.0 3.621
3.0 3.655
Perturbation-based 1.0 3.583
2.0 3.567
3.0 3.694
LibriSpeech-test-clean Model-based 1.0 6.81 7.222
2.0 6.927
3.0 6.927
Perturbation-based 1.0 6.965
2.0 6.874
3.0 6.684
@@ 4.5 Analysis of multi-epoch with different $r_{2}$
For the general tasks, following the experimental setup detailed in Section 4.1, $r_{2}$ coefficients were selected from the range [1, 1.5, 2, 3]. For vertically specific math tasks, using the configuration described in Section 4.2, $r_{2}$ coefficients were tested at values of [0.1, 0.2, 1, 2]. The corresponding experimental results are illustrated in Fig. 3(a) and Fig. 3(b).
@@
As shown in Fig. 3(a) and Fig. 3(b), our experimental results demonstrate two key conclusions: 1) Model performance varies on the leaderboard with different r2 coefficients. Thus, adjusting the r2 coefficient is necessary to prevent a rapid decline in rejection rate that causes overfitting. 2) The difficulty of learning varies across tasks. The r2 coefficient should be adjusted based on the rate of loss decrease for different tasks.